Setrakian Industries needs to raise $84.5 million to fund a new project. The company will sell bonds that have a coupon rate of 5.86 percent paid semiannually and that mature in 25 years. The bonds will be sold at an initial YTM of 6.58 percent and have a par value of $2,000. How many bonds must be sold to raise the necessary funds? (Round your intermediate calculations to two decimal places and final answer to the nearest whole number.)
Answer #1

Information provided:

Par value= future value= $2,000

Time= 25 years*2= 50 semi-annual periods

Coupon rate= 5.86%/2= 2.93%

Coupon payment= 0.0293*2,000= $58.60

Yield to maturity= 6.58%/2= 3.29%

The question is solved by first computing the price of the bond which is calculated with the help of present value.

The present value is computed by entering the below in a financial calculator:

FV= 2,000

N= 50

PMT= 58.60

I/Y= 3.29

Press the CPT key and PV to compute the present value.

The value obtained is 1,824.53.

Therefore, the price of the bond is $1,824.53.

Number of bonds to be raised= $84,500,000/ $1,824.53

                                                         = 46,313.2971

Hence, 46,313 bonds needs to be raised.

In case of any query, kindly comment on the solution.